# Persona Dynamics: Unveiling the Impact of Personality Traits on Agents in Text-Based Games

### v1

@@ Abstract
Seungwon Lim1       Seungbeen Lee2       Dongjun Min2       Youngjae Yu2,† Department of Artificial Intelligence, Yonsei University1    Department of Computer Science, Yonsei University2
@@ 2.3 Agent Implementation
where $Q(s_{t},a_{t}^{i})$ is the action value of i-th action among action candidates, optimized during training. $C(s_{t},a_{t}^{i}\mid p)\in\{-1,0,1\}$ represents the output of personality classifier. Given a pair of situation $s_{t}$ and action $a_{t}^{i}$, $-1$ denotes Low valence, 00 denotes Neutral valence, and $1$ denotes High valence regarding personality type $p$ to evaluate. The sign of $\gamma$ determines the direction of alignment. When $\gamma$ $>0$, it increases the Q-value of behaviors that match the personality trait, and vice versa. The absolute value determines the strength of the intended alignment. The agent’s action selection $a_{t}$ at state $s_{t}$ is determined by (2):
@@
where $|\mathcal{A}_{t}|$ denotes total number of action candidates.
@@
In this paper, we denote agents under the guidance of specific personality $p$ as $A_{p^{\uparrow}}$ when $\gamma$ $>0$ and $A_{p^{\downarrow}}$ when $\gamma$ $<0$ in equation (1), reflecting the valence of personality to guide. For example, an agent with high openness and low openness is denoted by $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$, respectively. Similarly, for a specific personality $p$, an action where the classifier predicted as High is denoted by $a_{p\text{\textuparrow}}$, and $a_{p\text{\textdownarrow}}$when Low.
@@ 3.1 Dataset Construction
Although these sentences collectively address various facets of each trait, there is a noticeable imbalance in their distribution, and approximately 70% of the items describe individuals exhibiting a high valence toward a given trait. To achieve balance, we systematically paraphrase these initial sentences to create 10 instances per trait (5 representing high valence and 5 representing low valence). Through this process, we obtain a total of 80 descriptions across 8 personality types Full examples are in 23 and 24.
@@
To make a single statement (e.g.‘I easily make new friends.’) into a detailed description (e.g. ‘I don’t worry about making new friends when moving schools’), we use GPT-4 to generate 300 diverse, common situations. These are divided into 30 subsets (e.g. Home and Family), each containing 10 scenarios (e.g. Kitchen, Garden). This approach, despite known biases in GPT-4 Gupta et al. (2023), helps data augmentation with diversity with minimal duplication West et al. (2021). Using 80 personality-describing sentences across 8 traits and the 300 situational seeds, we generate 5 sentences for each seed combination ($80\times 300\times 5=120,000$). The intermediate results of the dataset creation process and examples of generated samples are in Table 19 and 20.
@@ 3.2 Training and Performance
We trained Flan-T5 Chung et al. (2024) of various sizes on the dataset created in Section 3.1, and examined performance on personality classification using a diverse dataset which provided statements with annotated personality traits. Since BFI and SD-3 were used in dataset construction, we used IPIP Goldberg et al. (1999), another personality questionnaire, and Essays dataset Pennebaker and King (1999), containing author personality annotations across various types of writings as out-of-domain evaluation sets.
@@
Table 3: Game Scores on games of Jiminy Cricket. $A_{\text{N.P}}$(‘No Personality’) means no guidance with personality classifier, and the symbols (↑) and (↓) indicate high and low levels of each personality trait, respectively. Avg., Cnt. and Diff. are three criteria defined in Section 4.2. We only report 15 games here because in the remaining 10 games, agents of any personality type failed to score points in over 90 percent of attempts. Results for all games can be found in Table 12 and 13. For scores, bold indicates games satisfying the threshold condition for Cnt. The best scores are bolded and the second-best ones are underlined on metrics.
@@ 4.2 Results
$-\sum_{g}\text{if }[s(v,p)>s(\text{-})\text{ and }s(\bar{v},p)<s(\text{-})]$
@@ 4.3 Statistical analysis
Table 5: Analysis in last 50 Episodes based on each game agent’s movement trajectory in Zork1. Standard deviations and scores of omitted agents are provided in Table 16 and  17. The best scores are bolded and the second-best ones are underlined.
@@ 5.1 Trajectory Analysis
Figure 4 shows visual representations for each agent, with $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$ as representative examples. As shown in the Visit Count of $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$ in Table 5, both agents visited places near the starting point (Com.) during their respective training periods (8.96 and 8.02). However, while $A_{\text{Ope.}^{\downarrow}}$ rarely reaches places far from the starting point (Unc.), $A_{\text{Ope.}^{\uparrow}}$’s trajectory branches out in multiple directions. The visual example of all agent types can be found in Figure 13 to 16.
@@ 5.3 Walkthrough Analysis
Table 8: Analysis of personality traits in walkthrough actions for 25 games. Numbers (%) represent the ratio of each actions among all.
@@ 5.4 Learning Curve Analysis
Figure 5 shows the learning curves of $A_{\text{Ope.}^{\uparrow}}$, $A_{\text{N.P}}$, and $A_{\text{Ope.}^{\downarrow}}$ across 15 games to show how scores progress according to different learning stages. The area between the curves of $A_{\text{Ope.}^{\uparrow}}$and $A_{\text{Ope.}^{\downarrow}}$is notably large, which explains the score differences observed in Table 3.
@@ 6.1 Personality and LLMs
Assessing personality in advanced language models like GPT-4 and Claude has become an active research area recently  Miotto et al. (2022); Dorner et al. (2023). Most studies use psychometric tests originally designed for humans, like the Big Five Inventory, or machine-generated tests. However, these self-assessment tests lack detailed scenarios and are sensitive to factors like phrasing  Song et al. (2023); Caron and Srivastava (2023); Huang et al. (2023), making them unreliable for evaluating model personality.
To address these challenges, researchers are exploring alternative methodologies for more accurately assessing the personality traits of language models. One promising direction involves using interactive scenarios where the language model’s responses are evaluated by human judges or through automated sentiment analysis Gupta et al. (2024); Dorner et al. (2023); Frisch and Giulianelli (2024). This approach aims to capture more nuanced aspects of personality that may be overlooked by standard self-assessment tests.
@@ 9 Ethical Considerations
Misinterpreting personality-driven behaviors may lead users to form unrealistic expectations or emotional attachments to agents, potentially resulting in ethical concerns. To prevent such issues, we emphasize that personality traits in our agents are functional attributes aimed at improving alignment with human users, rather than indicators of sentient beings Safdar et al. (2020).
@@ Appendix A Overview
  • In Appendix B, we provide the detailed hyperparameters applied to train the DRRN agent used in our framework.
  • In Appendix C, we explain the detailed hyperparameters applied to train the personality classifier used in our framework.
  • In Appendix D, we provide examples of scenarios for games in Jiminy-Cricket benchmark.
@@
Appendix B DRRN Training Details
@@
Appendix C Personality classifier Training Details
@@
Appendix D Game Environments
D.1 Abbreviations
@@
D.2 Full Results on Table 3
Game scores and standard deviations across three different runs for all 25 games in the Jiminy-Cricket benchmark are presented in Table 12 and  13.
@@ D.2 Full Results on Table 3
D.3 Run Example of Jiminy Cricket Benchmark.
D.4 Learning Curve
We report the training progression through learning curves for all eight personality types, measured across a test suite of 15 games.
@@ D.4 Learning Curve
D.5 Comparison with Other Methodologies
@@
Average 1.7 4.8 4.6 4.6 4.9 4.1
Table 15: Comparison with previous text-game adventure agents. We report $A_{\text{Ope.}^{\uparrow}}$ as a representative example of the PANDA framework
D.6 Detailed Criteria for Place Classification

### v2

@@ Abstract
Seungwon Lim1     Seungbeen Lee2     Dongjun Min2     Youngjae Yu2 Department of Artificial Intelligence, Yonsei University1   Department of Computer Science, Yonsei University2
@@ 2.3 Agent Implementation
where $Q(s_{t},a_{t}^{i})$ is the action value of i-th action among action candidates, optimized during training. $C(s_{t},a_{t}^{i}\mid p)\in\{-1,0,1\}$ represents the output of personality classifier. Given a pair of situation $s_{t}$ and action $a_{t}^{i}$, 1-1- 1 denotes Low valence, 0 denotes Neutral valence, and 111 denotes High valence regarding personality type $p$ to evaluate. The sign of $\gamma$ determines the direction of alignment. When $\gamma$ >0>0> 0, it increases the Q-value of behaviors that match the personality trait, and vice versa. The absolute value determines the strength of the intended alignment. The agent’s action selection $a_{t}$ at state $s_{t}$ is determined by (2):
@@
where $|\mathcal{A}_{t}|$ denotes total number of action candidates. Qualitative examples of actions selected by agents with different personality traits are in Table 17.
@@
In this paper, we denote agents under the guidance of specific personality $p$ as $A_{p^{\uparrow}}$ when $\gamma$ >0>0> 0 and $A_{p^{\downarrow}}$ when $\gamma$ <0<0< 0 in equation (1), reflecting the valence of personality to guide. For example, an agent with high openness and low openness is denoted by $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$, respectively. Similarly, for a specific personality $p$, an action where the classifier predicted as High is denoted by $a_{p\text{\textuparrow}}$, and $a_{p\text{\textdownarrow}}$when Low.
@@ 3.1 Dataset Construction
Although these sentences collectively address various facets of each trait, there is a noticeable imbalance in their distribution, and approximately 70% of the items describe individuals exhibiting a high valence toward a given trait. To achieve balance, we systematically paraphrase these initial sentences to create 10 instances per trait (5 representing high valence and 5 representing low valence). Through this process, we obtain a total of 80 descriptions across 8 personality types Full examples are in 26 and 27.
@@
To make a single statement (e.g.‘I easily make new friends.’) into a detailed description (e.g. ‘I don’t worry about making new friends when moving schools’), we use GPT-4 to generate 300 diverse, common situations. These are divided into 30 subsets (e.g. Home and Family), each containing 10 scenarios (e.g. Kitchen, Garden). This approach, despite known biases in GPT-4 Gupta et al. (2023), helps data augmentation with diversity with minimal duplication West et al. (2021). Using 80 personality-describing sentences across 8 traits and the 300 situational seeds, we generate 5 sentences for each seed combination (80×300×5=120,00080$\times$ 300$\times$ 5=120,00080 × 300 × 5 = 120 , 000). The intermediate results of the dataset creation process and examples of generated samples are in Table  22 and 23.
@@ 3.2 Training and Performance
We trained Flan-T5 Chung et al. (2024) of various sizes on the dataset created in §3.1, and examined performance on personality classification using a diverse dataset which provided statements with annotated personality traits. Since BFI and SD-3 were used in dataset construction, we used IPIP Goldberg et al. (1999), another personality questionnaire, and Essays dataset Pennebaker and King (1999), containing author personality annotations across various types of writings as out-of-domain evaluation sets.
@@
Table 3: Game Scores on games of Jiminy Cricket. $A_{\text{N.P}}$(‘No Personality’) means no guidance with personality classifier, and the symbols (↑) and (↓) indicate high and low levels of each personality trait, respectively. Avg., Cnt. and Diff. are three criteria defined in §4.2. We only report 15 games here because in the remaining 10 games, agents of any personality type failed to score points in over 90 percent of attempts. Results for all games can be found in Table 14 and 15. For scores, bold indicates games satisfying the threshold condition for Cnt. The best scores are bolded and the second-best ones are underlined on metrics.
@@ 4.2 Results
$\sum_{g}\text{if }[s(v,p)>s(\text{-})\text{ and }s(\bar{v},p)<s(\text{-})]$
@@ 4.3 Statistical analysis
Table 5: Analysis in last 50 Episodes based on each game agent’s movement trajectory in Zork1. Standard deviations and scores of omitted agents are provided in Table 19 and  20. The best scores are bolded and the second-best ones are underlined.
@@ 5.1 Trajectory Analysis
Figure 4 shows visual representations for each agent, with $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$ as representative examples. As shown in the Visit Count of $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Ope.}^{\downarrow}}$ in Table 5, both agents visited places near the starting point (Com.) during their respective training periods (8.96 and 8.02). However, while $A_{\text{Ope.}^{\downarrow}}$ rarely reaches places far from the starting point (Unc.), $A_{\text{Ope.}^{\uparrow}}$’s trajectory branches out in multiple directions. The visual example of all agent types can be found in Figure 14 to 17.
@@ 5.3 Walkthrough Analysis
Table 8: Analysis of personality traits in walkthrough actions for 25 games. Numbers (%) represent the ratio of each actions among all types of personality.
@@ 5.4 Learning Curve Analysis
Figure 5 shows the learning curves of $A_{\text{Ope.}^{\uparrow}}$, $A_{\text{N.P}}$, and $A_{\text{Ope.}^{\downarrow}}$ across 15 games to demonstrate score progression during reinforcement learning. When compared with agents of different personality types, 1) $A_{\text{Ope.}^{\uparrow}}$learns to achieve higher scores than $A_{\text{N.P}}$in numerous games, and 2) exhibits substantial gap from $A_{\text{Ope.}^{\downarrow}}$’s learning curve throughout the learning process. Learning curves for agents with other personality types are in Figure 7 to 13.
@@
6 Additional Analysis
6.1 Impact of Psychopathy
After analyzing and comparing the trajectories of game agents with walkthrough trajectories, we discovered that in some games, antisocial behavior is necessary to achieve higher game scores. For example, in the game “Zork 1”, the player should kill the troll to pass through the troll room by performing actions such as “kill troll with sword”.
We hypothesized that among the dark triad personality types, agents with the Psychopathy trait would display different behavioral patterns in these situations. To test this, we filtered high-psychopathy actions from §5.3, then investigated the probability of those actions performed by game agents of each personality type among the generated action candidates  [^2] Since analyzing experimental trajectories retrospectively to analyze an agent’s behavior in a specific state involves very few samples and is probabilistic, we verified action value by feeding specific states and action candidates to the policy network after RL learning was completed..
According to Table 9, $A_{\text{Psy.}^{\uparrow}}$ is more likely to select actions with high psychopathy in situations that require such behavior during the walkthrough, exhibiting on average 14% higher frequency compared to other agents. However, Table 5 shows that $A_{\text{Psy.}^{\uparrow}}$ has difficulty exploring the game space compared to other agents (especially $A_{\text{Ope.}^{\uparrow}}$), therefore significant score improvement is not observed in Table 3.
Also, $A_{\text{Psy.}^{\uparrow}}$ is followed by $A_{\text{Mac.}^{\uparrow}}$ in terms of frequency of actions with high psychopathy in Table 9, which we attribute to the established correlation between high psychopathy and high machiavellianism traits McHoskey et al. (1998); Ali and Chamorro-Premuzic (2010).
Agent $P(a_{c}=a_{\text{Psy},\uparrow})$ Agent $P(a_{c}=a_{\text{Psy},\uparrow})$
$A_{\text{Ope.}^{\uparrow}}$ 12.1 $A_{\text{Ope.}^{\downarrow}}$ 14.4
$A_{\text{Con.}^{\uparrow}}$ 12.9 $A_{\text{Con.}^{\downarrow}}$ 12.9
$A_{\text{Ext.}^{\uparrow}}$ 18.1 $A_{\text{Ext.}^{\downarrow}}$ 1.1
$A_{\text{Agr.}^{\uparrow}}$ 9.8 $A_{\text{Agr.}^{\downarrow}}$ 18.4
$A_{\text{Neu.}^{\uparrow}}$ 10.8 $A_{\text{Neu.}^{\downarrow}}$ 9.2
$A_{\text{Psy.}^{\uparrow}}$ 25.5 $A_{\text{Psy.}^{\downarrow}}$ 1.4
$A_{\text{Mac.}^{\uparrow}}$ 20.7 $A_{\text{Mac.}^{\downarrow}}$ 13.0
$A_{\text{Nar.}^{\uparrow}}$ 13.6 $A_{\text{Nar.}^{\downarrow}}$ 2.3
Table 9: Analysis for actions demonstrating high psychopathy tendencies. With 10 action candidates generated by the action generator, $P(a_{c}=a_{\text{Psy},\uparrow})$ denotes the probability that $a_{\text{Psy},\uparrow}$ to be selected based on the softmax distribution of action values assigned by each agent’s trained policy network.
6.2 Alignment with Human Gameplay
To evaluate the alignment with human decision patterns across agents with different personality traits, we obtained human trajectories and investigated which personality type of agent shows the most similar action decision pattern to humans. For a given game state $s_{t}$ and action candidates, we measured the concordance between the action chosen by the human players and the actions chosen by agents with different personality types. For comparison, we also measured concordance with a DRRN Agent without any assigned personality traits ($A_{\text{N.P}}$).
As shown in Table 10, the behavior patterns of DRRN Agents without assigned personality traits ($A_{\text{N.P}}$) show high concordance with $A_{\text{Ope.}^{\uparrow}}$ and $A_{\text{Agr.}^{\uparrow}}$, while human players show high concordance with $A_{\text{Psy.}^{\uparrow}}$ and $A_{\text{Mac.}^{\uparrow}}$ compared to other personality types.
Agent Ope. Con. Ext. Agr. Neu. Psy. Mac. Nar.
Human $A_{p^{\uparrow}}$ 11.9 11.8 18.5 7.1 11.2 28.8 19.9 11.6
$A_{p^{\downarrow}}$ 9.7 10.8 1.4 11.4 6.1 1.4 11.0 1.6
Agent Ope. Con. Ext. Agr. Neu. Psy. Mac. Nar.
$A_{\text{N.P}}$ $A_{p^{\uparrow}}$ 18.8 10.4 9.2 14.8 11.5 9.6 12.0 12.1
$A_{p^{\downarrow}}$ 11.6 9.7 14.6 10.4 9.7 11.6 10.3 8.1
Table 10: Concordance rates between 16 distinct personality type agents and both human players and baseline DRRN agents, demonstrating differential behavior patterns across personality types.
6.3 Correlation between personality traits
In the real world, each personality does not exist independently in isolation but rather exists in complex interrelationship with other personality types Van der Linden et al. (2010). To investigate patterns of personality correlations in personality types defined over text-based game actions, we examined the interrelationships between personality annotations of actions chosen by each agent.
The personality pair showing the highest correlation is openness and extraversion (0.92), which is also confirmed by demonstrating the longest Trajectory Length in Table 5. Meanwhile, the five personality types of the Big Five and the three personality types of the Dark Triad exhibit minimal inter-dimensional correlation. Notably, psychopathy demonstrates a strong negative correlation with conscientiousness (-0.71). This suggests a trade-off between antisocial and conscientious behaviors, which may serve as a key consideration for agent design.
Figure 6: Correlation for each personality type. Correlation values are based on the agreement of personality annotations of actions that constitute trajectories across all games. Actions with neutral valence were not considered.
@@ 7.1 Personality and LLMs
Assessing personality in advanced language models like ChatGPT and Claude has become an active research area recently  Miotto et al. (2022); Dorner et al. (2023). Most studies use psychometric tests originally designed for humans, like the Big Five Inventory, or machine-generated tests. However, these self-assessment tests lack detailed scenarios and are sensitive to factors like phrasing  Song et al. (2023); Caron and Srivastava (2023); Huang et al. (2023), making them unreliable for evaluating model personality.
To address these challenges, researchers are exploring alternative methodologies for more accurately assessing the personality traits of language models. One promising direction involves using interactive scenarios where the language model’s responses are evaluated by human judges or through automated sentiment analysis Gupta et al. (2024); Dorner et al. (2023); Frisch and Giulianelli (2024); Lee et al. (2024). This approach aims to capture more nuanced aspects of personality that may be overlooked by standard self-assessment tests.
@@ 10 Ethical Considerations
Misinterpreting personality-driven behaviors may lead users to form unrealistic expectations or emotional attachments to agents, potentially resulting in ethical concerns davidson2020human. To prevent such issues, we emphasize that personality traits in our agents are functional attributes aimed at improving alignment with human users, rather than indicators of sentient beings Safdar et al. (2020).
@@ Appendix A Overview
  • In Appendix B, we provide the detailed hyperparameters applied to train the DRRN agent and personality classifier used in our framework .
  • In Appendix C, we provide information and various examples of games of Jiminy Cricket benchmark.
  • In Appendix D, we present the full results of results in main page, along with additional materials and analyses that were not covered on the main page.
@@
Appendix B Training Details
B.1 Details for training DRRN Agent
@@
B.2 Details for training Personality classifier
@@
Appendix C Games in Jiminy Cricket environment
C.1 Abbreviations
@@
C.2 Example of Jiminy Cricket Benchmark.
C.3 World Visualization and Trajectory of Agent
We applied the visualization method presented in §5.1 to all agents, with results shown in Figure 14 to 17. Additionally, we visualize the map of the game Zork-1, Zork-2, and Zork-3 from Jiminy Cricket benchmark in in Figure 19 to 21.
Appendix D Full Results & Additional Analysis
D.1 Full Results on Table 3
Game scores and standard deviations across three different runs for all 25 games in the Jiminy-Cricket benchmark are presented in Table 14 and  15.
@@ D.1 Full Results on Table 3
D.2 Learning Curve
We report the training progression through learning curves for all eight personality types, measured across a test suite of 15 games in Figure 7 to 13.
@@ D.2 Learning Curve
D.3 Comparison with Other Methodologies
@@
D.4 Qualitative examples of action decisions from agents
In Table 17, we list examples of actions annotated with personality types from randomly selected game situations.
Personality Game Observation & Action
High Openness Seastalker [Ob] You’re in the pilot’s seat of the SCIMITAR … The power reactor is off and open. [Act] put capsule in reactor
Low Openness Lurkinghorror [Ob] The darkness before you, now visible, is a creature. … You can feel the smooth stone vibrating in your hand. [Act] wait
High Conscientiousness Zork1 [Ob] This appears to have been an artist’s studio. … Loosely attached to a wall is a small piece of paper. [Act] read paper
Low Conscientiousness Wishbringer [Ob] You’re in a dark underground chamber. Cold, smelly tunnels wander away to the south and east. … [Act] drop all except stone
High Extraversion Suspect [Ob] Ballroom, South Entrance. Samuel Ostmann, enjoying himself hugely as a vampire, is here. … After a short pause, the band begins to play "Tennessee Waltz". [Act] ask ostmann about himself
Low Extraversion Ballyhoo [Ob] Inside Prop Tent. This dimly lighted tent houses a ragtag collection of outrageous props and wacky devices, including a midget automobile that’s garaged in here. … [Act] hide
High Agreeableness Lurkinghorror [Ob] A really whiz-bang pc is right inside the door. … Sitting at a terminal is a hacker whom you recognize. [Act] give carton to hacker
Low Agreeableness Ballyhoo You step up into the trailer. … Sitting on an upper bunk, with his arms folded and legs swinging, is Comrade Thumb. [Act] close door
High Neuroticism Witness [Ob] The body of Mr. Linder is still crumpled in a heap on the floor. [Act] hide behind lounge
Low Neuroticism Ballyhoo [Ob] You nonchalantly walk away from the long line. The monkey on your back idly grooms areas of your scalp. [Act] get in long line
High Psychopathy Zork1 [Ob] This is a small room with passages to the east and south and a forbidding hole leading west. … The flat of the troll’s axe skins across your forearm. [Act] kill troll with sword
Low Psychopathy Plunderedhearts [Ob] The large cupboard is not flush against the doorsill. … A large cupboard against the aft wall blocks the doorway to the captain’s quarters. [Act] open door
High Machiavellianism Seastalker [Ob] But Commander Bly has heard your order… [Act] ask bly about private matter
Low Machiavellianism Suspect [Ob] Ostmann would love to buy the farm, but we turned him down. … Colonel Marston stops here. Linda stops here. [Act] ask about himself
High Narcissism Seastalker [Ob] You are standing in a small room, that of the protective goddess Selkis. … A beautiful ruby cluster, sparkling like fire, sits on the floor. [Act] get ruby
Low Narcissism Infidel [Ob] You emerge into the warm night air of summer. … A midget decked out in a Russian general’s uniform is standing before the drinking fountain. [Act] help midget
Table 17: Qualitative examples of observation ([Ob]) & action ([Act]) pairs annotated with 16 personality types.
@@ D.4 Qualitative examples of action decisions from agents
Average 0.9 3.5 3.4 3.6 3.8 4.1
Table 18: Comparison with previous text-game adventure agents. We report $A_{\text{Ope.}^{\uparrow}}$ as a representative example of the PANDA framework.
D.5 Detailed Criteria for Place Classification